OpenStudy (anonymous):

at what point do y = 2/3 x and y = (cos^3x)^1/2 intersect?

OpenStudy (anonymous):

So I was trying to solve this problem, lemme put it in a nicer format \[y = \frac{ 2 }{ 3 }x\] \[y = \sqrt{\cos ^{3}x}\]

OpenStudy (anonymous):

but when I put them equal to each other \[\frac{ 4 }{9 }x ^{2} = \cos ^{3}x\] But I can't simplfy it any further to solve for x.... or can I? This section I can use the calculator... but IDK how to solve for x.

OpenStudy (loser66):

hope this help.

OpenStudy (anonymous):

Well how do I determine what the point is?? Can I do it mechanically do must I use a calculator?